OpenStudy (marialupossalazar):

If 3/4 and 12/x are equivalent, what is the value of x?

OpenStudy (zzr0ck3r):

3*? = 12

OpenStudy (marialupossalazar):

so whats the answer?

OpenStudy (zzr0ck3r):

3 times what equals 12?

OpenStudy (marialupossalazar):

4

OpenStudy (zzr0ck3r):

ok so multiply 4 by 4 and what do you get?

OpenStudy (marialupossalazar):

16

OpenStudy (zzr0ck3r):

that is the answer we had \(\dfrac{3}{4}\) and we want to make that equal to \(\dfrac{12}{x}\) Since we needed to multiply the top by \(4\) to get \(12\), that means we must multiply the bottom by \(4\) to get the answer.

OpenStudy (zzr0ck3r):

in other words \(\dfrac{3}{4}=\dfrac{3}{4}\times 1=\dfrac{3}{4}\times \dfrac{4}{4}=\dfrac{12}{16}\)

OpenStudy (marialupossalazar):

oh okay thank you

OpenStudy (zzr0ck3r):

np